-- Main.jill - 15 Jan 2007

-- Main.gredel - 17 Jan 2007

work on rotator all day, possible cause identified, command rotator --> spinning motor shaft found, but gear not moving telescope handover 5:50 pm, start lbccontrol

00:57 start 25 x bias to check --> looks OK aborted after 13 bias frames

01:04 sky flats U-band dither 5 NOTE OB SkyFlatU_5 (and B, V, etc.) DO NOT SEEM TO DITHER!! create own OBs for skyflats with dither pattern 01:07 problem with telescope, TCS error, restart, cannot open doors lbccontrol hanging init lbccontrol, connect to LBCT hangs interface

log analyser 577 2007/01/17 01:29:57.361000 CAMERA CCDCTRL detectors temperature: 173.4 [Kelvin] 03:3576 2007/01/17 01:29:55.435369 received >UISTATUS< command 575 2007/01/17 01:29:53.623627 HKEEPING TEMPERAT temperature: 127.4 [Kelvin] 574 2007/01/17 01:29:53.623552 HKEEPING VACUUM pressure: 2.78E-06 [mbar] 573 2007/01/17 01:29:46.593593 message >command rejected!< sent to the user UI 572 2007/01/17 01:29:46.593490 unknown command >TCSON< 571 2007/01/17 01:29:46.593460 received >TCSON< command 570 2007/01/17 01:29:46.593110 TCS IIF_Authorize completed

* TCS Problems * John Little & Michelle (via phone) working to get building doors responding again to commands,

01:56 doors open, preset to L95 --> OK load Bfastextra --> lbccontrol gui very very slow, seems to hang, recovers after several minutes 02:04 IA images moved from /home/telescope/Repository to /Repository at the beginning of night, lbcfpia fails. cannot read image produced by Bfastextra FITS_READ_ERORR: No extension for given extname, extver, and/or extlevel found

02:17 check, take 5x bias --> WHERE ARE THE IMAGES??? they are neither in /home/telescope/Repository nor in /Repository THE LAST 5 BIASES ARE NOT IN THE ARCHIVE EITHER

02:28 reboot mcu, take 5 bias again --> where are the images

managed to contact Riccardo Smareglia on his way down, logged on from basecamp, working on archive right now

--> all moved back to /home/telescope/Repository, possibly permission issues which caused the failures --> never do last-minute changes just before you leave the mountain

03:29 back on sky; check, 5 x bias --> looks good

03:37 image analysis; mirror at -7 C, ambient -9


Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
4020 0 0 0 0 0
  1. -130 225 14 -241 -253
  2. -117 313 44 -21 -1

03:53 L95_B,V 1 and 2 sec in B and V each --> 1.1 arcsec

03:56 L95_U 1, 2, 10 sec --> 1.2 arcsec

READY TO GO

03:59 Trilling/KBO1 Target F1a --> 0.86 arcsec in V-band start with OB100a, then 100b, 101a, 101b, etc., following cadence requirements

04:24 image analysis


Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-1119 -668 -200 167 304 -229 127 6 -2 -62 180 -17

04:33 Trilling OB102a etc. --> seeing 0.7 in V night is not photometric

seeing up to 1.1 arcsec in V

05:02 IA
Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
247 -200 -130 -192 69 87

05:07 Trilling OB 105a etc.

humidity climbing, around 85% images good sub-arcsec, continues w/o image analysis

05:57 humidity 92% close shutters

Topic revision: r5 - 17 Jan 2007, gredel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback